# Sectorial covers over fanifolds

### v2

@@ 1. Introduction
The Fukaya category of a symplectic manifold is a highly nontrivial invariant. It is an $A_{\infty}$-category whose homotopy category has Lagrangians as objects and morphisms between two Lagrangians is the linear span of their intersection points up to suitable perturbation. The $A_{\infty}$-structure is controlled by counts of pseudoholomorphic disks with Lagrangian boundary conditions, which depend on the global geometry of the symplectic manifold. Usually, computing Fukaya categories is extremely hard because of this global nature. However, under suitable assumptions, Fukaya categories exhibit good local-to-global behaviors which allows us to reduce complicated analysis on the global geometry of symplectic manifolds to more tractable one.
For instance, via sectorial descent recently established in [GPS2] by Ganatra–Pardon–Shende, the wrapped Fukaya category $\mathcal{W}(X)$ of a Weinstein sector $X$ can be computed by gluing that of local pieces, provided its Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$. More precisely, the canonical functor from the homotopy colimit
$\displaystyle\operatorname{colim}_{\emptyset\neq I\subset\{1,\ldots,n\}}\mathcal{W}(\bigcap_{i\in I}X_{i})\to\mathcal{W}(X)$
@@
induced by the pushforward functors along inclusions of Weinstein sectors from [GPS1] is a pretriangulated equivalence. In other words, wrapped Fukaya categories are cosheaves with respect to Weinstein sectorial covers. The result is valid also for partially wrapped Fukaya categories of stopped Weinstein sectors with mostly Legendrian stops away from boundaries.
@@
Considering its computational power, given a Weinstein sector, it is natural to seek its Weinstein sectorial cover. Moreover, when working with homological mirror symmetry(HMS), one often needs such a cover to be compatible with that of its mirror. For instance, if the latter cover corresponds to an open cover of a relative skeleton, then a lift of it to a Weinstein sectorial cover will have the desired compatibility. So far, there are not so many interesting examples of Weinstein sectorial covers [CKL, GJ, GL22]. One reason for this might be easier access to other local-to-global behaviors [GS22, GS23, Lee16, MSZ, PS1, PS2, PS3], which requires us to compute at least locally a skeleton. Although they might work to solve a given problem, construction of Weinstein sectorial covers still should be a fundamental geometric question.
@@
In this paper, we give a Weinstein sectorial cover of the stopped Weinstein sector associated with a fanifold introduced in [GS23] by Gammage–Shende to discuss HMS for more general spaces than toric stacks on the $B$-side. It is a stratified space obtained by gluing rational polyhedral fans of cones, which provides the organizing topological and discrete data for HMS at large volume. To a fanifold $\Phi$ they associated an algebraic space $\mathbf{T}(\Phi)$ obtained as the gluing of the toric varieties $T_{\Sigma}$ associated with the fans $\Sigma$. Based on an idea from SYZ fibrations, they constructed its mirror $\widetilde{\mathbf{W}}(\Phi)$ by inductive Weinstein handle attachments. The guiding principle behind their construction of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ was that the FLTZ Lagrangians $\mathbb{L}(\Sigma)$ and the projections $\mathbb{L}(\Sigma)\to\Sigma$ should glue to yield its relative skeleton $\widetilde{\mathbb{L}}(\Phi)$ and a map $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$.
As predicted by Gammage–Shende [GS23, Remark 4.5], the canonical lifts of the projections $\mathbb{L}(\Sigma)\to\Sigma$ glue to yield a version of $A$-side SYZ fibration $\widetilde{\mathbf{W}}(\Phi)\to\Phi$, after some modifications.
Theorem 1.1 ([Mor, Theorem 1.2]).
There is a filtered stratified fibration $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ restricting to $\pi$, which is homotopic to a filtered stratified integrable system with noncompact fibers. If the fan $\Sigma_{S}$ associated with any stratum $S\subset\Phi$ is proper, then the homotopy becomes trivial.
The lift $\bar{\pi}$ completes a global combinatorial duality for the mirror pair $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ over $\Phi$. Moreover, it enables us to lift any cover of $\Phi$ to $\widetilde{\mathbf{W}}(\Phi)$. Inspired by the idea in [GPS2, Example 1.34] and using this aspect of $\bar{\pi}$, after some modification, we construct a Weinstein sectorial cover of $\widetilde{\mathbf{W}}(\Phi)$ with the desired compatibility in Section $4$.
@@
When $\Phi$ is the fanifold from [GS23, Example 4.23], $\widetilde{\mathbb{L}}(\Phi)$ coincides with the skeleton $\operatorname{Core}(H)$ [GS22, Theorem 6.2.4] of a very affine hypersurface $H$. In Section $5$, we show that our cover gives a lift of the open cover [GS22, Corollary 4.3.2] of $\operatorname{Core}(H)$, giving an affirmative answer to the conjecture raised in [GS22, Section 2.4] and making the discussion of [GS22, Section 2.1] completely rigorous.
Via sectorial descent we obtain a description of HMS for $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ as an isomorphism of cosheaves of categories. See Section $4.3$ for the precise definitions of the associated cosheaves. The isomorphism can be regarded as a dual to [GS23, Theorem 5.3] and encodes the gluing of coherent constructible correspondence [Kuw20] for local pieces.
@@
There is an isomorphism
$\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$
of cosheaves of categories over $\Phi$ whose global section yields an equivalence
$\operatorname{Fuk}(\widetilde{\mathbf{W}}(\Phi))\simeq\operatorname{Coh}(\mathbf{T}(\Phi)).$
Our construction of the cover is close to the more general strategy suggested in [BC23] by Bai–Côté. When $X=T^{*}M$, they triangulate $M$ and cover $X$ by the cotangent bundles $T^{*}\operatorname{star}(v_{\alpha})$ of the stars of the vertices $v_{\alpha}$. When $X$ is a polarizable Weinstein manifold, they deform $X$ to another Weinstein manifold with arborial skeleton which by [BC23, Section 4.2] admits a canonical Whitney stratification, in particular, a triangulation due to Goresky. Cover the deformation of $X$ by arboreal thickenings of the stars of the vertices. Unfortunately, there are obstructions to making their strategy rigorous which involves developing a good theory of arboreal sectors.
@@ Acknowledgements.
The author was supported by SISSA PhD scholarships in Mathematics. This work was partially supported by JSPS KAKENHI Grant Number JP23KJ0341.
2. Gluing of stopped Weinstein sectors
In this section, we review some basic properties of Liouville sectors introduced in [GPS1, Section 2] by Ganatra–Pardon–Shende and how to glue them together. Throughout the paper, we work with Liouville manifolds of finite type so that any Liouville manifold $X$ will be the completion of some Liouville domain $X_{0}$. Then its skeleton $\operatorname{Core}(X)$ will always be compact. We denote by $\partial_{\infty}X_{0}$ the actual boundary of $X_{0}$, which can be identified with the ideal boundary $\partial_{\infty}X$ of $X$ to preserve the symbols $\partial X_{0},\partial X$.
2.1. Stopped Weinstein sectors
All Liouville manifolds of our interest are Weinstein. Namely, the Liouville vector field $Z$ of each $X$ is gradient-like for a Morse–Bott function $\phi\colon X\to\mathbb{R}$ which is constant on the cylindrical end $(\mathbb{R}_{t\geq 0}\times\partial_{\infty}X,e^{t}(\lambda|_{\partial_{\infty}X}))$. Hence $\operatorname{Core}(X)$ will also be isotropic by [CE12, Lemma 11.13(a)].
Definition 2.1 ([GPS1, Definition 2.4]).
A Liouville sector is a Liouville manifold-with-boundary satisfying the following equivalent conditions:
  • For some $\alpha>0$ there exists an $\alpha$-defining function $I\colon\partial X\to\mathbb{R}$ with $ZI=\alpha I$ near infinity and $dI|_{C}>0$ for the characteristic foliation of the hypersurface $\partial X\subset X$, where $C$ is oriented so that $\omega(C,N)>0$ for any inward pointing vector $N$.
  • The boundary of the contact hypersurface $\partial_{\infty}X\subset X$ is convex and there is a diffeomorphism $\partial X\cong F\times\mathbb{R}$ sending $C$ to the foliation of $F\times\mathbb{R}$ by leaves $\{pt\}\times\mathbb{R}$. We call the symplectic reduction $F$, which is a Liouville manifold, the symplectic boundary of $X$.
An inclusion $i\colon X\hookrightarrow X^{\prime}$ of Liouville sectors is a proper map which is a diffeomorphism onto its image, satisfying $i^{*}\lambda^{\prime}=\lambda+df$ for compactly supported $f$. A trivial inclusion of Liouville sectors is one for which $i(X)$ may be deformed into $X^{\prime}$ through Liouville sectors included in $X$.
Remark 2.2.
The inequality $dI|_{C}>0$ is equivalent to the Hamiltonian vector field $X_{I}$ being outward pointing along $\partial X$. Small deformations of $X$ within the class of Liouville manifolds-with-boundary are again Liouville sectors.
Definition 2.3 ([GPS1, Remark 2.8]).
An open Liouville sector is a pair $(X,\partial_{\infty}X)$ of an exact symplectic manifold $X$ and a contact manifold $\partial_{\infty}X$ together with a germ near +$\infty$+ ∞ of codimension 00 embedding of the symplectization $S\partial_{\infty}X$ into $X$. Here, the embedding strictly respects Liouville forms and the pair $(X,\partial_{\infty}X)$ is exhausted by Liouville sectors. Being exhausted by Liouville sectors means that any subset of $X$, which away from a compact subset of $X$ is the cone over a compact subset of $\partial_{\infty}X$, is contained in some Liouville sector $X_{0}\subset X$ with $\partial_{\infty}X_{0}\subset\partial_{\infty}X$.
Example 2.4.
For any manifold $M$ the cotangent bundle $T^{*}M$ is an open Liouville sector. An exhaustion is given by the family of Liouville subsectors $T^{*}M_{0}$ for compact codimension 00 submanifolds-with-boundary $M_{0}\subset M$. For any Liouville sector $X$ its interior $X^{\circ}=X\setminus\partial X$ is an open Liouville sector.
For a Liouville sector $X$ we denote by $\mathcal{W}(X)$ the wrapped Fukaya category [GPS1, Definition 3.36]. It is an $A_{\infty}$-category whose objects are exact cylindrical at infinity Lagrangians inside $X$ and whose morphisms are calculated by wrapping Lagrangians. According to [GPS1, Section 3.6], any inclusion $X\hookrightarrow X^{\prime}$ of Liouville sectors induces the pushforward functor $\mathcal{W}(X)\to\mathcal{W}(X^{\prime})$ which makes the assignment $X\mapsto\mathcal{W}(X)$ covariantly functorial. We use the same symbols to denote the counterparts of open Liouville sectors.
Wrapped Fukaya categories are invariant under deformations in the following sense.
Lemma 2.5 ([GPS2, Lemma 3.4]).
Along a trivial inclusion $X\hookrightarrow X^{\prime}$ of Liouville sectors, the pushforward functor defines a canonical quasi-equivalence $\mathcal{W}(X)=\mathcal{W}(X^{\prime})$.
Definition 2.6 ([GPS2, Section 1.1]).
A stop $\mathfrak{f}$ of a Liouville sector $X$ is a closed subset of $(\partial_{\infty}X)^{\circ}=\partial_{\infty}X\setminus\partial\partial_{\infty}X$. We call the pair $(X,\mathfrak{f})$ a stopped Liouville sector. An inclusion $(X,\mathfrak{f})\hookrightarrow(X^{\prime},\mathfrak{f}^{\prime})$ of stopped Liouville sectors is an inclusion of Liouville sectors $X\hookrightarrow X^{\prime}$ with $\mathfrak{f}^{\prime}\cap(\partial_{\infty}X)^{\circ}\subset\mathfrak{f}$.
When $\mathfrak{f}=\emptyset$, the above definition specializes to Definition 2.1. For a stopped Liouville sector $(X,\mathfrak{f})$ we denote by $\mathcal{W}(X,\mathfrak{f})$ the partially wrapped Fukaya category [GPS2, definition 2.31]. It is an $A_{\infty}$-category whose objects are exact cylindrical at infinity Lagrangians inside $X$ disjoint at infinity from $\mathfrak{f}$ and whose morphisms are calculated by wrapping Lagrangians in $(\partial_{\infty}X)^{\circ}\setminus\mathfrak{f}$. Similar to wrapped Fukaya categories, any inclusion $(X,\mathfrak{f})\hookrightarrow(X^{\prime},\mathfrak{f}^{\prime})$ of stopped Liouville sectors induces the pushforward functor $\mathcal{W}(X,\mathfrak{f})\to\mathcal{W}(X^{\prime},\mathfrak{f}^{\prime})$ which makes the assignment $(X,\mathfrak{f})\mapsto\mathcal{W}(X,\mathfrak{f})$ covariantly functorial.
Although the partially wrapped Fukaya category is defined for an arbitrary stopped Liouville sector $(X,\mathfrak{f})$, the results in [GPS2] are generally sharpest when $X$ is Weinstein and $\mathfrak{f}$ is mostly Legendrian.
Definition 2.7 ([GPS2, Definition 1.7]).
A closed subset $\mathfrak{f}$ of a contact manifold $Y^{2n-1}$ is mostly Legendrian if it admits a decomposition $\mathfrak{f}=\mathfrak{f}^{\text{subcrit}}\cup\mathfrak{f}^{\text{crit}}$ for which $\mathfrak{f}^{\text{subcrit}}$ is closed and is contained in the smooth image of a second countable manifold of dimension $<n-1$, and $\mathfrak{f}^{\text{crit}}\subset Y\setminus\mathfrak{f}^{\text{subcrit}}$ is a Legendrian submanifold. The notion of a mostly Lagrangian closed subset of a symplectic manifold is defined analogously.
2.2. Weinstein pairs
Lemma 2.8 ([GPS1, Lemma 2.13]).
Let $\bar{X}_{0}$ be a Liouville domain and $A\subset\partial_{\infty}\bar{X}_{0}$ a codimension 00 submanifold-with-boundary for which there exists a function $I\colon A\to\mathbb{R}$ with $R_{\lambda}I>0$ such that the contact vector field $V_{I}$ is outward pointing along $\partial A$. Then $X=\bar{X}\setminus(A^{\circ}\times\mathbb{R}_{>0})$ is a Liouvile sector where $\bar{X}$ denotes the conic completion of $\bar{X}_{0}$.
Definition 2.9 ([GPS1, Definition 2.14]).
A sutured Liouville domain $(\bar{X}_{0},F_{0})$ is a Liouville domain $\bar{X}_{0}$ together with a codimension $1$ submanifold-with-boundary $F_{0}\subset\partial_{\infty}\bar{X}_{0}$ and a contact form $\lambda$ defined over a neighborhood $\operatorname{Nbd}_{\partial_{\infty}\bar{X}_{0}}(F_{0})$ of $F_{0}$ in $\partial_{\infty}\bar{X}_{0}$ such that $(F_{0},\lambda|_{F_{0}})$ is a Liouville domain. A sutured Liouville manifold is a Liouville manifold $\bar{X}$ together with a codimension $1$ submanifold-with-boundary $F_{0}\subset\partial_{\infty}\bar{X}$ and a contact form $\lambda$ defined over a neighborhood $\operatorname{Nbd}_{\partial_{\infty}\bar{X}}(F_{0})$ of $F_{0}$ in $\partial_{\infty}\bar{X}$ such that $(F_{0},\lambda|_{F_{0}})$ is a Liouville domain.
Remark 2.10.
Any sutured Liouville domain $(\bar{X}_{0},F_{0})$ defines a Weinstein pair in the sense of [Eli, Section 2] if $\lambda$ extends to $\bar{X}_{0}$ and there exists a function $\phi\colon\bar{X}_{0}\to\mathbb{R}$ which makes the triples $(\bar{X}_{0},\lambda,\phi)$ and $(F_{0},\lambda|_{F_{0}},\phi|_{F_{0}})$ Weinstein.
Definition 2.11 ([GPS1, Remark 2.16]).
Given a sutured Liouville domain $(\bar{X}_{0},F_{0})$ let $\mathbb{L}_{0}\subset\bar{X}_{0}$ be the locus of points which do not escape to the complement of $\operatorname{Core}(F_{0})$ at infinity. We call $\mathbb{L}_{0}$ the relative skeleton of $(\bar{X}_{0},F_{0})$ and denote it by $\operatorname{Core}(\bar{X}_{0},F_{0})$. Analogous terminology applies to sutured Liouville manifolds.
Remark 2.12.
The locus $\mathbb{L}\subset\bar{X}$ is necessarily noncompact unless $F_{0}$ is empty. When both $\bar{X},F_{0}$ are Weinstein, $\mathbb{L}$ is a singular isotropic spine. In particular, $\bar{X}$ deforms down to a small regular neighborhood of $\mathbb{L}$.
Given a sutured Liouville domain $(\bar{X}_{0},F_{0})$, the Reeb vector field $R_{\lambda}$ is transverse to $F_{0}$ as $d\lambda|_{F_{0}}$ is symplectic and determines a local coordinate charts $F_{0}\times\mathbb{R}_{|t|\leq\epsilon}\hookrightarrow\partial_{\infty}\bar{X}_{0}$ in which we have $\lambda=\lambda|_{F_{0}}+dt$. The contact vector field $V_{t}$ associated with the function $t$ is given by $Z|_{F_{0}}+t\partial_{t}$ which is outward pointing along $\partial(F_{0}\times\mathbb{R}_{|t|\leq\epsilon})$. Hence $(\bar{X}_{0},F_{0})$ determines a codimension 00 submanifold-with-boundary $A=F_{0}\times\mathbb{R}_{|t|\leq\epsilon}$ satisfying the hypotheses of Lemma 2.8. In particular, $(\bar{X}_{0},F_{0})$ gives rise to a Liouville sector.
Conversely, any Liouville sector arises, up to deformation, from a unique in the homotopical sense sutured Liouville domain. The following lemma plays a role to prove Lemma 2.17 below.
Lemma 2.13 ([GPS1, Lemma 2.23]).
Any pair $(\bar{X}_{0},A)$ satisfying the hypotheses of Lemma 2.8 arises, up to deformation, from a unique in the homotopical sense sutured Liouville domain.
Lemma 2.14 ([GPS1, Proposition 2.25]).
Let $X$ be a Liouville sector. Equip $\mathbb{C}$ the standard symplectic form $\omega_{\mathbb{C}}=dx\wedge dy$. Let $\lambda^{\alpha}_{\mathbb{C}}$ be the Liouville form associated with a Liouville vector field $Z^{\alpha}_{\mathbb{C}}=(1-\alpha)x\partial_{x}+\alpha y\partial_{y}$. Then any $\alpha$-defining function $I\colon\operatorname{Nbd}^{Z}_{X}(\partial X)\to\mathbb{R}$ extends to a unique identification
$(X,\lambda)=(F\times\mathbb{C}_{\operatorname{Re}\geq 0},\lambda_{F}+\lambda^{\alpha}_{\mathbb{C}}+df)$
valid over some neighborhoods of the respective boundaries, where $I=y$ is the imaginary part of the coordinate on $\mathbb{C}$, $(F,\lambda_{F})$ is a Liouville manifold and $f\colon F\times\mathbb{C}_{\operatorname{Re}\geq 0}\to\mathbb{R}$ satisfies the following conditions:
  • $f$ is supported inside $F_{0}\times\mathbb{C}$ for some Liouville domain $F_{0}\subset F$.
  • $f$ coincides with some $f_{\pm\infty}\colon F\to\mathbb{R}$ for sufficiently large $|I|$.
Definition 2.15 ([GPS1, Section 2.7]).
Let $X$ be a Liouville sector. The convex completion is the pair $(\bar{X},\lambda_{\bar{X}})$ of an exact symplectic manifold $\bar{X}=X\cup_{\operatorname{Nbd}\partial(F\times\mathbb{C}_{\operatorname{Re}\geq 0})}\operatorname{Nbd}(F\times\mathbb{C}_{\operatorname{Re}\leq 0})$ and a $1$-form $\lambda_{\bar{X}}=\lambda_{F}+\lambda^{1/2}_{\mathbb{C}}+d\tilde{f}$ for an extension $\tilde{f}\colon\operatorname{Nbd}(F\times\mathbb{C}_{\operatorname{Re}\leq 0})\to\mathbb{R}$ of $f\colon\operatorname{Nbd}\partial(F\times\mathbb{C}_{\operatorname{Re}\geq 0})\to\mathbb{R}$.
Lemma 2.16 ([GPS1, Lemma 2.31]).
If the extension $\tilde{f}\colon F\times\mathbb{C}_{\operatorname{Re}<\epsilon}\to\mathbb{R}$ is bounded uniformly in $C^{\infty}$ and satisfies the two bulleted properties from Lemma 2.14, then the convex completion $(\bar{X},\lambda_{\bar{X}})$ defines a Liouville manifold.
Lemma 2.17 ([GPS1, Lemma 2.32]).
Any Liouville sector arises, up to deformation, from a unique in the homotopical sense sutured Liouville manifold. Moreover, the convex completion of the Liouville sector $X$ associated with a sutured Liouville manifold $(\bar{X},F_{0})$ coincides with $\bar{X}$, and the inclusion $X\hookrightarrow\bar{X}$ is the obvious one.
2.3. Gluing of Weinstein pairs
Given a Weinstein pair $(\bar{X}_{0},F_{0})$, by [Eli, Proposition 2.9] one can always modify the Liouville form $\lambda$ for $\omega$ and the Morse–Bott function $\phi_{F_{0}}$ to be adjusted in the following sense.
Definition 2.18 ([Eli, Section 2]).
Let $(\bar{X}_{0},F_{0})$ be a Weinstein pair, $\phi_{F_{0}}\colon F_{0}\to\mathbb{R}$ a Morse–Bott function for which $Z_{F_{0}}=(\omega|_{F_{0}})^{\#}\lambda|_{F_{0}}$ is gradient-like and $U_{\epsilon}(F_{0})\subset\partial_{\infty}\bar{X}_{0}$ the contact surrounding of $F_{0}$. We call a pair $(\lambda_{0},\phi_{0})$ of a Liouville form $\lambda_{0}$ for $\omega$ and a smooth function $\phi_{0}\colon\bar{X}_{0}\to\mathbb{R}$ adjusted to $(\bar{X}_{0},F_{0})$ if
  • $Z_{0}=\omega^{\#}\lambda_{0}$ is tangent to $\partial_{\infty}\bar{X}_{0}$ on $U_{\epsilon}(F_{0})$ and transverse to $\partial_{\infty}\bar{X}_{0}$ elsewhere;
  • $Z_{0}|_{U_{\epsilon}(F_{0})}=Z_{F_{0}}+u\partial_{u}$;
  • the attractor $\bigcap_{t>0}Z^{-t}_{0}(\bar{X}_{0})$ coincides with $\operatorname{Core}(\bar{X}_{0},F_{0})$;
  • the function $\phi_{0}$ is Morse–Bott for which $Z_{0}=(\omega)^{\#}\lambda_{0}$ is gradient-like satisfying $\phi_{0}|_{U_{\epsilon}(F_{0})}=\phi_{F_{0}}+\frac{1}{2}u^{2}$ and whose critical values are not more than $\phi_{0}|_{\partial U_{\epsilon}(F_{0})}$.
Definition 2.19 ([Eli, Section 3.1]).
Let $(\bar{X}_{0},\lambda,\phi)$ be a Weinstein domain. A splitting for $(\bar{X}_{0},\lambda,\phi)$ is a hypersurface $(P,\partial_{\infty}P)\subset(\bar{X}_{0},\partial_{\infty}\bar{X}_{0})$ satisfying the following conditions:
  • $\partial_{\infty}P$ and $P$ respectively split $\partial_{\infty}\bar{X}_{0}$ and $\bar{X}_{0}$ into two parts $\partial_{\infty}\bar{X}_{0}=Y_{-}\cup Y_{+}$ with $\partial_{\infty}Y_{-}=\partial_{\infty}Y_{+}=Y_{-}\cap Y_{+}=\partial_{\infty}P$ and $\bar{X}_{0}=\bar{X}_{0-}\cup\bar{X}_{0+}$ with $\partial\bar{X}_{0-}=P\cup Y_{-},\partial\bar{X}_{0+}=P\cup Y_{+},\bar{X}_{0-}\cap\bar{X}_{0+}=P$.
  • The Liouville vector field $Z=\omega^{\#}\lambda$ is tangent to $P$.
  • There exists a hypersurface $(F_{0},\partial_{\infty}F_{0})\subset(P,\partial_{\infty}P)$ which is Weinstein for $\lambda|_{F_{0}}$, tangent to $Z$ and intersects transversely all leaves of the characteristic foliation of $P$.
The above hypersurface $(F_{0},\partial_{\infty}F_{0})$ is called the Weinstein soul for the splitting hypersurface $P$. Due to Lemma 2.21 below, $P$ is contactomorphic to the contact surrounding $U_{\epsilon}(F_{0})$ of its Weinstein soul.
Definition 2.20 ([Eli, Section 2]).
Let $F_{0}$ be a closed hypersurface in a $(2n-1)$-dimensional manifold and $\xi$ a germ of a contact structure along $F_{0}$ which admits a transverse contact vector field $V$. The invariant extension of the germ $\xi$ is the canonical extension $\hat{\xi}$ on $F_{0}\times\mathbb{R}$, which is invariant with respect to translations along the second factor and whose germ along any slice $F_{0}\times\{t\},t\in\mathbb{R}$ is isomorphic to $\xi$.
Lemma 2.21 ([Eli, Lemma 2.6]).
Let $F_{0}$ be a closed $(2n-2)$-dimensional manifold and $\xi$ a contact structure on $P=F_{0}\times[0,\infty)$ which admits a contact vector field $V$ inward transverse to $F_{0}\times\{0\}$ such that its trajectories intersecting $F_{0}\times\{0\}$ fill the whole manifold $P$. Then $(P,\xi)$ is contactomorphic to $(F_{0}\times[0,\infty),\hat{\xi})$, where $\hat{\xi}$ is the invariant extension of the germ of $\xi$ along $F_{0}\times\{0\}$. Moreover, for any compact set $C\subset P$ with $F_{0}\times\{0\}\subset\operatorname{Int}C$ there exists a contactomorphism $(P,\xi)\to(F_{0}\times[0,\infty),\hat{\xi})$, which equals the identity on $F_{0}\times\{0\}$ and sends $V|_{C}$ to the vector field $\partial_{t}$.
Definition 2.22 ([Eli, Section 3.1]).
Let $(\bar{X}_{0},F_{0}),(\bar{X}^{\prime}_{0},F^{\prime}_{0})$ be Weinstein pairs with adjusted Liouville forms and Morse–Bott functions $(\lambda,\phi),(\lambda^{\prime},\phi^{\prime})$. Suppose that there is an isomorphism $(F_{0},\lambda|_{F_{0}},\phi|_{F_{0}})\cong(F^{\prime}_{0},\lambda^{\prime}|_{F^{\prime}_{0}},\phi^{\prime}|_{F^{\prime}_{0}})$ of Weinstein manifolds. Extend it to their contact surroundings $U_{\epsilon}(F_{0})\cong U_{\epsilon}(F^{\prime}_{0})$. Then the gluing of $(\bar{X}_{0},F_{0}),(\bar{X}^{\prime}_{0},F^{\prime}_{0})$ is given by
$\bar{X}_{0}\#_{F_{0}\cong F^{\prime}_{0}}\bar{X}^{\prime}_{0}=\bar{X}_{0}\cup_{U_{\epsilon}(F_{0})\cong U_{\epsilon}(F^{\prime}_{0})}\bar{X}^{\prime}_{0}$
with glued Liouville form and Morse–Bott function.
By definition we have
$\operatorname{Core}(\bar{X}_{0}\#_{F_{0}\cong F^{\prime}_{0}}\bar{X}^{\prime}_{0})=\operatorname{Core}(\bar{X}_{0},F_{0})\cup_{\operatorname{Core}(F_{0})\cong\operatorname{Core}(F^{\prime}_{0})}\operatorname{Core}(\bar{X}^{\prime}_{0},F^{\prime}_{0}).$
3. Large volume limit fibrations over fanifolds
In this section, after the construction of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ associated with a fanifold $\Phi$ introduced in [GS23] by Gammage–Shende, we review that of the filtered stratified fibration $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ from Theorem 1.1 restricting to that $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$ of the relative skeleton $\widetilde{\mathbb{L}}(\Phi)$. The stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ is obtained by inductively attaching products of the cotangent bundles of real tori and strata of $\Phi$. Each step requires us to modify Weinstein structures near gluing regions.

### v3

@@ 1. Introduction
Sectorial descent established in [GPS2] by Ganatra–Pardon–Shende is one of the local-to-global behaviors of Fukaya categories. It enables us to compute the wrapped Fukaya category $\mathcal{W}(X)$ of a Weinstein sector $X$ by gluing that of local pieces, provided its Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$. In other words, wrapped Fukaya categories are cosheaves with respect to Weinstein sectorial covers. The result is valid also for partially wrapped Fukaya categories of stopped Weinstein sectors with mostly Legendrian stops away from boundaries.
@@
Considering its computational power, given a Weinstein sector, it is natural to seek its Weinstein sectorial cover. Moreover, when working with homological mirror symmetry(HMS), one often needs such a cover to be compatible with that of its mirror. So far, there are not so many interesting examples of Weinstein sectorial covers [CKL, GJ, GL22]. One reason for this might be easier access to other local-to-global behaviors [GS22, GS23, Lee16, MSZ, PS1, PS2, PS3]. Although they might work to solve a given problem, construction of Weinstein sectorial covers still should be a fundamental geometric question.
@@
In this paper, we give a Weinstein sectorial cover of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ associated with a fanifold $\Phi$. A fanifold is introduced in [GS23] by Gammage–Shende to discuss HMS for more general spaces than toric stacks on the $B$-side. It is a stratified space obtained by gluing rational polyhedral fans of cones, which provides the organizing topological and discrete data for HMS at large volume. To $\Phi$ they associated an algebraic space $\mathbf{T}(\Phi)$ obtained as the gluing of the toric varieties $T_{\Sigma}$ for the fans $\Sigma$. Based on an idea from SYZ fibrations, they constructed its mirror $\widetilde{\mathbf{W}}(\Phi)$ by inductive Weinstein handle attachments.
The guiding principle behind their construction of $\widetilde{\mathbf{W}}(\Phi)$ was that the FLTZ Lagrangians $\mathbb{L}(\Sigma)$ and the projections $\mathbb{L}(\Sigma)\to\Sigma$ should glue to yield its relative skeleton $\widetilde{\mathbb{L}}(\Phi)$ and a map $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$. As predicted in [GS23, Remark 4.5], the canonical lifts of the projections $\mathbb{L}(\Sigma)\to\Sigma$ glue to yield a version of $A$-side SYZ fibration $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ after some modifications [Mor, Theorem 1.2]. The map $\bar{\pi}$ enables us to lift any cover of $\Phi$ to $\widetilde{\mathbf{W}}(\Phi)$. Inspired by the idea in [GPS2, Example 1.34] and using this aspect of $\bar{\pi}$, after some modification, we construct a Weinstein sectorial cover of $\widetilde{\mathbf{W}}(\Phi)$ with the desired compatibility in Section $3$.
@@
When $\Phi$ is the fanifold from [GS23, Example 4.23], $\widetilde{\mathbb{L}}(\Phi)$ coincides with the skeleton $\operatorname{Core}(H)$ [GS22, Theorem 6.12] of a very affine hypersurface $H$. In Section $4$, we show that our cover gives a lift of the open cover [GS22, Corollary 4.8] of $\operatorname{Core}(H)$, giving an affirmative answer to the conjecture raised in [GS22, Section 2.5] and making the heuristic of [GS22, Section 2.1] completely rigorous.
Via sectorial descent we obtain a description of HMS for $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ as an isomorphism of cosheaves of categories. See Section $3.4$ for the precise definitions of the associated cosheaves. The isomorphism can be regarded as a dual to [GS23, Theorem 5.3] and encodes the gluing of coherent constructible correspondence [Kuw20] for local pieces.
@@
There is an isomorphism $\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$ of cosheaves of categories over $\Phi$ whose global section yields an equivalence $\operatorname{Fuk}(\widetilde{\mathbf{W}}(\Phi))\simeq\operatorname{Coh}(\mathbf{T}(\Phi))$.
Our construction of the cover is close to the more general strategy suggested in [BC23] by Bai–Côté. When $X=T^{*}M$, they triangulate $M$ and cover $X$ by the cotangent bundles $T^{*}\operatorname{star}(v_{\alpha})$ of the stars of the vertices $v_{\alpha}$. When $X$ is a polarizable Weinstein manifold, they deform $X$ to another Weinstein manifold with arboreal skeleton which by [BC23, Section 4.2] admits a canonical Whitney stratification, in particular, a triangulation due to Goresky. Cover the deformation of $X$ by arboreal thickenings of the stars of the vertices. Unfortunately, there are obstructions to making their strategy rigorous which involves developing a good theory of arboreal sectors.
@@ Acknowledgements.
The author was supported by SISSA PhD scholarships in Mathematics. He is grateful to an anonymous referee for careful reading and helpful comments. This work was partially supported by JSPS KAKENHI Grant Number JP23KJ0341.
2. Large volume limit fibrations
In this section, after the construction of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ associated with a fanifold $\Phi$ introduced in [GS23] by Gammage–Shende, we review that of the map $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ from [Mor, Theorem 1.2] restricting to the map $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$ from the relative skeleton $\widetilde{\mathbb{L}}(\Phi)$. The stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ is obtained by inductively attaching products of the cotangent bundles of real tori and strata of $\Phi$. Each step requires us to modify Weinstein structures near gluing regions.